OpenStudy (anonymous):

find the slope if it exists x+8y=16

OpenStudy (saifoo.khan):

solve the equation for y = mx + b Or in other words, make y the subject. then compare the values. the value with x, will be the value for slope.

OpenStudy (anonymous):

i get confused with these i know that y=mx+b but it is all confusing to me will ya help me solve this please

OpenStudy (saifoo.khan):

umm, ALright..

OpenStudy (saifoo.khan):

as you have to make "y" the subject, so: x+8y=16 8y = x + 16 divide 8 from both sides, y = x + 16 ------ 8 \[y = \frac 18x + 2\]

OpenStudy (jamesj):

@saifoo, you've made a sign error

OpenStudy (jamesj):

If x + 8y = 16 then 8y = -x + 16 So y = (-1/8)x + 2

OpenStudy (saifoo.khan):

as you have to make "y" the subject, so: x+8y=16 8y = -x + 16 divide 8 from both sides, y = -x + 16 ------ 8 \[y= -\frac18x+2\]

OpenStudy (saifoo.khan):

thanks man

OpenStudy (jamesj):

sure thing